Three friends shared a Pizza, and each person ate their share at different times in the afternoon. If friend A came first and ate 1/3 of the pizza, and friend B ate 1/2 of what was left by friend A, how much was left for friend C?
A) 2/3
B) 1/6
C) 1/3
D) 3/4

Answer by rfer(16322)

1/3+(2/3)(1/2)=
1/3+1/3=2/3
3/3-2/3=1/3 left